# Data-Juicer 2.0: Cloud-Scale Adaptive Data Processing for Foundation Models

### v1

@@ 6.2.2. Automatic GPU Resource Allocation
As the results show, using GPUs saves at least 50% of processing time for all selected model-based OPs, especially for large, slow models like BLIP-2 (Li et al., 2023a) used in image_captioning_mapper and SDXL model (Rombach et al., 2022) in image_diffusion_mapper. Compared to the CPU version, GPU usage and adaptive resource allocation are much more efficient and necessary. Due to the GPU memory limit of around 80GB, we can only allocate at most 4 models of image_captioning_mapper OP on a single GPU, so the number of processors for this OP is automatically reduced to 32, preventing OOM errors.
6.2.3. Batched Data Processing
@@
Figure 9. Data processing efficiency of OP-wise hierarchical parallelism. “bs” denotes the batch size, and “nt” dennotes the number of threads.
6.2.4. Automatic OP-wise Hierarchical Parallelism
To verify the efficiency of the automatic OP-wise hierarchical parallel strategy, we select the multimodal OP image_aspect_ratio_filter as example. We conduct experiments with various processor counts, batch sizes, and thread numbers, covering three parallel levels: multiprocessing, batched processing, and multithreading. This OP processes approximately 560k image-text pairs. Fig. 9b demonstrates the time consumption and efficiency gains.
@@ 6.2.4. Automatic OP-wise Hierarchical Parallelism
7. Conclusions
Data-Juicer 2.0 emerges as a versatile scaffold in the evolution of foundation models, providing efficient and adaptive data processing solutions for handling the diversity and scale of modern datasets. Our re-envisioned architecture leverages multi-layered adaptability to coordinate different modules, operators, and runtime environments. Extensive evaluations reveal its high performance across diverse cloud-scale workloads. By open-sourcing this system, we aim to foster a vibrant community of contributors and users, encouraging collaborative development and driving innovation to underpin the next generation of foundation models. Future works include further enhancing the system in terms of adaptability using model-driven agents (Gao et al., 2024), scalability by optimizing the single-node transmission bottleneck in Ray’s head node (Team, 2024d), and efficiency by supporting GPU backend engines like NeMo Curator (Jennings et al., [n.d.]).

### v2

@@ H.2.2 Automatic GPU resource allocation
As the results show, using GPUs saves at least 50% of processing time for all selected model-based OPs, especially for large, slow models like BLIP-2 [37] used in image_captioning_mapper and SDXL model [56] in image_diffusion_mapper. Compared to the CPU version, GPU usage and adaptive resource allocation are much more efficient and necessary. Due to the GPU memory limit of around 80GB, we can only allocate at most 4 models of image_captioning_mapper OP on a single GPU, so the number of processors for this OP is automatically reduced to 32, preventing OOM errors.
H.2.3 Batched data processing
@@
Figure 10: Data processing efficiency of OP-wise hierarchical parallelism. “bs” denotes the batch size, and “nt” dennotes the number of threads.
H.2.4 Automatic OP-wise hierarchical parallelism
To verify the efficiency of the automatic OP-wise hierarchical parallel strategy, we select the multimodal OP image_aspect_ratio_filter as an example. We conduct experiments with various processor counts, batch sizes, and thread numbers, covering three parallel levels: multiprocessing, batched processing, and multithreading. This OP processes approximately 560k image-text pairs. Fig. 10b demonstrates the time consumption and efficiency gains.